OpenStudy (anonymous):

find the slant height of a cone with surface area 381.77 ft and radius 6.2 ft ill become your fan and give you an award

OpenStudy (campbell_st):

well the formula is \[S.A. = \pi r s + \pi r^2\] r = radius and s = slant height so substituting your known information \[381.77 = \pi \times 6.2 \times s + \pi \times 6.2^2\] just solve the equation for s

OpenStudy (anonymous):

so S= 241.54

OpenStudy (campbell_st):

well to start you need to subtract the area of a circle then divide by 6.2 pi to find s \[s= \frac{381.77 - \pi \times 6.2^2}{\pi \times 6.2}\]

OpenStudy (anonymous):

ok so 343.33/6.2 times pie

OpenStudy (campbell_st):

no do the calculation correctly \[s =\frac{(381.77 - 120.76)}{19.48}\]

OpenStudy (anonymous):

that equal 13.39